This year Jamie is three times as old as her brother john. Next year she will be twice as old, how old is jamie ?
Answer by vksarvepalli(154)

let the present age of john be x
so the age of jamie this year will be 3x
next year age of john will be = x+1
and age of jamie will be 3x+1
given Next year she will be twice as old as john.
so, 3x+1 = 2(x+1)
=> 3x +1 = 2x+2
=> x=1 and 3x will be = 3

so, age of jamie is 3 years.
Ans: 3 years